OpenStudy (anonymous):

Explain how you would find the exact value of 3 over sqrt 8.

OpenStudy (lgbasallote):

using rationalization \[\huge \frac{3}{\sqrt 8} \implies \frac{3}{\sqrt 8} \times \frac{\sqrt 8}{\sqrt 8}\] does that help?

OpenStudy (precal):